### v2

# MOREL: Enhancing Adversarial Robustness through Multi-Objective Representation Learning

Abstract
Extensive research has shown that deep neural networks (DNNs) are vulnerable to slight adversarial perturbations—small changes to the input data that appear insignificant but cause the model to produce drastically different outputs. In addition to augmenting training data with adversarial examples generated from a specific attack method, most of the current defense strategies necessitate modifying the original model architecture components to improve robustness or performing test-time data purification to handle adversarial attacks. In this work, we demonstrate that strong feature representation learning during training can significantly enhance the original model’s robustness. We propose MOREL, a multi-objective feature representation learning approach, encouraging classification models to produce similar features for inputs within the same class, despite perturbations. Our training method involves an embedding space where cosine similarity loss and multi-positive contrastive loss are used to align natural and adversarial features from the model encoder and ensure tight clustering. Concurrently, the classifier is motivated to achieve accurate predictions. Through extensive experiments, we demonstrate that our approach significantly enhances the robustness of DNNs against white-box and black-box adversarial attacks, outperforming other methods that similarly require no architectural changes or test-time data purification. Our code is available at https://github.com/salomonhotegni/MOREL.
1 Introduction
The deployment of deep neural networks (DNNs) in critical vision applications such as autonomous driving and medical diagnosis (Bojarski et al., 2016; Miotto et al., 2018) underscores the need for robust models capable of reliably handling real-world scenarios. However, extensive research has demonstrated that DNNs are vulnerable to adversarial examples—inputs crafted by adding imperceptible perturbations that can cause the model to make incorrect predictions with high confidence (Nguyen et al., 2015; Szegedy et al., 2013; Goodfellow et al., 2014). This vulnerability poses significant challenges to the security and reliability of AI systems, especially in safety-critical environments. To mitigate the risks posed by adversarial attacks, various defense strategies have been proposed. A common approach is adversarial training (Madry et al., 2017), where models are trained on adversarial examples generated from a specific attack method. In addition, to improve robustness, most existing defenses require modifications to the original model architecture (Panousis et al., 2021; Liu et al., 2024; Mohammed et al., 2024; Zhou et al., 2023), introducing additional complexity and often being architecture-dependent. Some approaches also involve test-time data purification (Meng & Chen, 2017; Song et al., 2017; Cohen & Giryes, 2024; Tang & Zhang, 2024), which increases latency, limiting their practical applicability.
In this paper, we propose a novel method named Multi-Objective REpresentation Learning (MOREL) that addresses these challenges by focusing on robust feature representation learning. MOREL encourages the model to produce consistent features for inputs within the same class, despite adversarial perturbations. By enhancing the robustness of feature representations, MOREL strengthens the model’s inherent ability to differentiate between classes, making it more resilient to adversarial attacks. The core of our approach is a multi-objective optimization framework that simultaneously optimizes two key objectives: enhancing adversarial robustness and maintaining high classification accuracy. We achieve this by embedding natural and adversarial features into a lower-dimensional space, where cosine similarity and contrastive loss functions are applied to align and tightly cluster these features. The classifier is concurrently motivated to achieve accurate predictions. This approach ensures that the model’s learned representations are robust to adversarial perturbations while preserving the information necessary for accurate classification. The embedding space used during training is discarded, allowing the model to retain its original structure and computational efficiency during inference. This characteristic distinguishes MOREL from many existing defense strategies that either involve architectural changes or rely on additional modules during inference. Through extensive experiments (Sec. 4), we demonstrate that our approach significantly enhances the robustness of DNN models against white-box and black-box adversarial attacks, outperforming existing adversarial training methods that similarly require no architectural changes or test-time data purification, in terms of the accuracy-robustness trade-off. In summary, our key contributions are:
  • We propose Multi-Objective REpresentation Learning (MOREL), a framework that enhances the robustness of deep neural networks by aligning natural and adversarial features in a shared embedding space during training while preserving the model’s original structure for practical deployment.
  • We approach the challenge of improving adversarial robustness and maintaining high accuracy as a multi-objective optimization task, effectively balancing these objectives to enhance the accuracy-robustness trade-off.
  • We demonstrate through extensive experiments that models trained with MOREL outperform those trained with existing adversarial training methods, supporting our hypothesis that strong feature representation learning enhances model robustness.
2 Related Work
2.1 Adversarial Training
Adversarial training, introduced by Madry et al. (2017), has emerged as one of the most effective defenses against adversarial attacks. The core idea involves augmenting the training data with adversarial examples generated using methods like Projected Gradient Descent (PGD). While standard adversarial training has proven effective against known attacks, it often results in models becoming overly specialized to the specific types of adversarial examples used during training (Tsipras et al., 2018). To address this limitation, several variants of adversarial training have been proposed. Kannan et al. (2018) introduced Adversarial Logit Pairing (ALP), which enhances robustness by pairing logits from adversarial and clean examples during training. Building on this, they proposed Clean Logit Pairing (CLP), which further refines the approach by focusing specifically on randomly selected clean training examples. Ding et al. (2018) advanced the field with Max-Margin Adversarial (MMA) training, which pushes decision boundaries further from data points, thereby offering enhanced robustness. The TRADES method by Zhang et al. (2019) marked a significant leap forward by explicitly balancing the trade-off between robustness and accuracy through a regularized loss function that minimizes the Kullback-Leibler divergence between predictions on natural and adversarial examples. This was further refined by MART Wang et al. (2019), which focuses on the robustness of misclassified examples, addressing vulnerabilities near the decision boundary. Despite these advancements, common limitations persist, including the challenge of maintaining strong robustness while achieving high accuracy on clean data. Building on these state-of-the-art adversarial training methods, our approach, MOREL, addresses these challenges by strengthening the robustness of DNNs through a robust feature representation learning technique. By considering a multi-objective optimization framework, MOREL aims to achieve the best possible trade-offs between robustness and accuracy—an aspect that, to our knowledge, has not been fully explored in previous work.
2.2 Insights from Contrastive Learning
To enhance the learning of robust features in the context of adversarial training, our method also draws insights from recent advances in contrastive learning. Contrastive learning has been shown to be effective in producing robust and well-structured feature representations by encouraging similar samples to be closer in the embedding space while pushing dissimilar samples apart (Chen et al., 2020; Gidaris et al., 2018; He et al., 2020). Specifically, Khosla et al. (2020) extend the principles of contrastive learning to a supervised setting. This method leverages label information to group similar examples (i.e., those sharing the same class label) closer together in the feature space. This work informs the design of our embedding space in MOREL, where we apply a multi-positive contrastive loss function (Khosla et al., 2020; Tian et al., 2024) to align natural and adversarial features. By doing so, MOREL not only enhances robustness against adversarial attacks but also ensures that the learned features are tightly clustered and well-separated across different classes, improving both robustness and accuracy.
3 Methods
We consider a supervised classification problem where a DNN model $f$ parameterized by $\theta\in\Omega$ learns to map an input image $x\in\mathbb{R}^{d}$ to a target class $f(x)=y\in\{1,...,c\}$ where $c\in\mathbb{N}$. An adversarial example $x^{\prime}\in\mathbb{R}^{d}$ is an image obtained by adding a small perturbation of at most $\epsilon>0$ to every pixel of $x$ such that $f(x)\neq f(x^{\prime})$. With a given $l_{p}$-based adversarial region $\mathcal{R}_{p}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{p}\leq\epsilon\}$, the aim of adversarial training (Madry et al., 2017) is typically to approximately minimize the risk on the data distribution $\mathcal{D}$ over adversarial examples:
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\max_{x^{\prime}\in\mathcal{R}_{p}(x,\epsilon)}\mathcal{L}(\theta,f(x^{\prime}),y)\right]$ (1)
where $\mathcal{L}$ is the loss function. The approximate solutions to the inner maximization problem are derived using a specific attack method to generate adversarial examples, while the outer minimization problem involves training on these generated examples.
3.1 Adversarial Attacks
To generate adversarial examples for training, we use the Projected Gradient Descent (PGD) attack (Madry et al., 2017). It is an iterative method that generates adversarial examples by iteratively applying small perturbations to the input. Given an input image $x$, the true label $y$, a loss function $\mathcal{L}(\theta,x,y)$, and a model parameterized by $\theta$, the PGD attack generates an adversarial example $x^{\prime}$ through the following iterative process for a predefined number of iterations:
${x^{\prime}}^{0}=x$ (2)
${x^{\prime}}^{i+1}=\text{Proj}_{\mathcal{R}(x,\epsilon)}\left({x^{\prime}}^{i}+\eta\cdot\text{sign}\left(\nabla_{x}\mathcal{L}(\theta,{x^{\prime}}^{i},y)\right)\right)$ (3)
where, ${x^{\prime}}^{i}$ is the adversarial example at the $i$-th iteration, $\eta$ the step size, $\epsilon$ the maximum perturbation allowed, and $\text{Proj}_{\mathcal{R}(x,\epsilon)}$ the projection operator that ensures the adversarial example remains within the $\epsilon$-ball centered at $x$. Especially, we consider the $l_{\infty}$-based adversarial region:
$\mathcal{R}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{\infty}\leq\epsilon\}.$
3.2 Multi-Objective Representation Learning
Training a robust model often results in a decrease in test accuracy. The goal of adversarial robustness is then to mitigate the trade-off between accuracy and robustness, thereby enhancing the model’s performance on both natural and adversarial examples (Zhang et al., 2019; Raghunathan et al., 2020). We approach this challenge as a multi-objective optimization problem. The first objective is to constrain the model to produce features that are as similar as possible for input images within the same class, and as dissimilar as possible from feature distributions of other classes, despite perturbations. The second objective is to enhance the model’s accuracy. We denote the model encoder as $g$ (typically the model without its final layer) and the classifier as $h$ (typically the final layer). Let =$\mathcal{B}=\big{\{}x_{i}\in\mathbb{R}^{d}\mid i\in\{1,...,n\}\big{\}}$ be a batch of $n$ natural images with classes $\big{\{}y_{i}\in\{1,...,c\}\mid i\in\{1,...,n\}\big{\}}=\mathcal{Y}$, and =$\mathcal{B}^{\prime}=\big{\{}x_{i}^{\prime}\in\mathcal{R}_{p}(x_{i},\epsilon)\mid i\in\{1,...,n\}\big{\}}$ its adversarial batch. The encoder then produces features [^1] We use the matrix notation $(z_{i})_{i=1}^{n}=Z\in\mathbb{R}^{n\times o}$, where $Z$ is the concatenation of the $n$ vectors $z_{i}$, each of dimension $o$. In the remainder of this paper, when we use the $\bigoplus$ operation over these matrices, it refers to a concatenation operation.:
$g(\mathcal{B})=(z_{i})_{i=1}^{n},\ \text{and}\ g(\mathcal{B}^{\prime})=(z^{\prime}_{i})_{i=1}^{n}.$ (4)
3.2.1 Embedding Space with Class-Adaptive Multi-Head Attention
During training, we consider an embedding space that includes a linear layer $L_{\text{e}}$ to project the features from the encoder into a lower-dimensional space:
$L_{\text{e}}((z_{i})_{i=1}^{n})=(s_{i})_{i=1}^{n},\ \text{and}\ L_{\text{e}}((z^{\prime}_{i})_{i=1}^{n})=(s^{\prime}_{i})_{i=1}^{n}.$ (5)
The lower-dimensional features are then grouped according to their classes:
$(s_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}(s_{i}^{y})_{i=1}^{n_{y}},\ \text{and}\ (s^{\prime}_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}.$ (6)
where $n_{y}$ is the number of features of class $y$ present.
Additionally, a class-adaptive multi-head attention module $M_{\text{e}}$ enables interaction within each lower-dimensional feature group, resulting in richer feature representations. This module functions similarly to the multi-head attention mechanism in the vision transformer (Dosovitskiy et al., 2020; Xiong et al., 2020), where the linearly embedded image patches can be viewed as a lower-dimensional feature group. The key distinction is that our multi-head attention module operates on features from different images (instead of features from the patches of the same image), and we omit any positional embedding mechanism since the position of a feature within its lower-dimensional feature group is irrelevant in our case (otherwise, this would imply keeping track of the position of an image within its batch).
More precisely, given a lower-dimensional feature group $(s_{i}^{y})_{i=1}^{n_{y}}\ \left(\text{or}\ ({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$, the module $M_{\text{e}}$ produces the final embedded feature group $(t_{i}^{y})_{i=1}^{n_{y}}\ \left(\text{or}\ ({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$ via Algorithm 1.
Algorithm 1 Class-Adaptive Multi-Head Attention
0:  A feature group $(s_{i}^{y})_{i=1}^{n_{y}}=S_{y}\in\mathbb{R}^{n_{y}\times b}$, from class $y$.
0:  The availability of learnable triplet weight matrices $W_{j}^{Q}$, $W_{j}^{K}$ and $W_{j}^{V}\in\mathbb{R}^{b\times b_{j}}\ (b_{j}=b/m)$ for each head $j\in\{1,...,m\}$ as well as an additional learnable weight matrix $W^{O}\in\mathbb{R}^{mb_{j}\times b}$.
1:  for $j=1$ to $m$ do
2:     Normalize $S_{y}$ via layer normalization.
3:     Project $S_{y}$ through linear transformations:$Q_{j}=S_{y}W_{j}^{Q},\ K_{j}=S_{y}W_{j}^{K},\ \text{and}\ V_{j}=S_{y}W_{j}^{V}.$
4:     Get the attention score: $A_{j}={\tt softmax}\left(\dfrac{Q_{j}K_{j}^{T}}{\sqrt{b_{j}}}\right)$
5:     Compute the $j^{\text{th}}$ head output $O_{j}=A_{j}V_{j}$.
6:  end for
7:  Concatenate the outputs from the $m$ heads and project the result through a linear transformation:$O={\tt concat}(O_{1},\dots,O_{m})W^{O}$
8:  return  $(t_{i}^{y})_{i=1}^{n_{y}}=S_{y}+O$
This approach takes advantage of the global context understanding property of the attention mechanism (Dosovitskiy et al., 2020; Han et al., 2022) to capture dependencies and relationships across features within the same group (class). During model evaluation on the test set, the embedding space is discarded, keeping the original model architecture unchanged.
Figure 1: Overview of our proposed MOREL method. During training, encoder $g(\cdot)$ features are projected into a lower-dimensional space using a linear layer $L_{\text{e}}$. A class-adaptive multi-head attention module $M_{\text{e}}$ then facilitates interaction within each of the obtained feature groups. The natural and adversarial embedded features are aligned using $\mathcal{L}_{cosine}$ (8), and $\mathcal{L}_{csl}$ (9) ensures tight clustering of the class-informed embedded features. Simultaneously, the classifier $h(\cdot)$ is encouraged to make accurate predictions. During evaluation, the embedding space is discarded, preserving the original model architecture.
3.2.2 Multi-Objective Optimization
In multi-objective optimization, the goal is to simultaneously optimize two or more conflicting objectives, which requires balancing trade-offs to find solutions that satisfy all objectives (losses) to an acceptable degree (Hotegni et al., 2024; Coello, 2007; Marler & Arora, 2004). We define the loss function for robustness based on the $l_{2}-$normalized batch features from the embedding space:
$\left\|\bigoplus\limits_{y\in\{1,...,c\}}(t_{i}^{y})_{i=1}^{n_{y}}\right\|_{2}=(t_{i})_{i=1}^{n},\ \text{and}\ \left\|\bigoplus\limits_{y\in\{1,...,c\}}({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right\|_{2}=(t^{\prime}_{i})_{i=1}^{n}.$ (7)
Cosine Similarity Loss:
The cosine similarity loss function measures the cosine similarity between pairs of feature vectors, encouraging the model to produce similar features for a natural image and its adversarial example in the embedding space. It is calculated as follows, considering $L_{\text{e}}$ outputs:
$\mathcal{L}_{\text{cosine}}=1-\dfrac{1}{n}\sum_{i=1}^{n}\frac{t_{i}\cdot t^{\prime}_{i}}{\|t_{i}\|\|t^{\prime}_{i}\|}$ (8)
[…]
Multi-Positive Contrastive Loss:
The multi-positive contrastive loss function (Khosla et al., 2020; Tian et al., 2024) encourages the model to bring the features of the same class closer while pushing the features of different classes apart, considering the natural features from $M_{\text{e}}$:
$\mathcal{L}_{csl}=\sum_{j\in\{1,...,2n\}}\dfrac{-1}{|\mathcal{P}(j)|}\sum_{p\in\mathcal{P}(j)}\text{log}\dfrac{\text{exp}(t_{j}\cdot t_{p}/\tau)}{\sum_{q\in\mathcal{Q}(j)}\text{exp}(t_{j}\cdot t_{q}/\tau)}$ (9)
where $\tau\in\mathbb{R}^{+}$ is a scalar parameter, $\mathcal{Q}(j)=\{1,...,2n\}\setminus\{j\}$ and $\mathcal{P}(j)=\{p\in\mathcal{Q}(j)\mid y_{p}=y_{j}\}$ with $y_{p}$ and $y_{j}$ the class labels of $t_{p}$ and $t_{j}$.
The loss function for robustness is then defined as follows:
$\mathcal{L}_{1}=\mathcal{L}_{\text{cosine}}+\alpha\mathcal{L}_{csl}$ (10)
with $10<\alpha<$ 1.
To improve accuracy, we recommend using loss functions that induce robustness in the classifier $h(\cdot)$, such as TRADES (Zhang et al., 2019) or MART (Wang et al., 2019). TRADES is defined as a Cross-Entropy loss ($\mathcal{L}_{\text{ce}}$) regularized by the Kullback-Leibler divergence ($\mathcal{D}_{\text{kl}}$) between the model’s predictions on natural and adversarial examples:
$\mathcal{L}_{2}=\mathcal{L}_{\text{ce}}(h((z_{i})_{i=1}^{n}),\mathcal{Y})+\frac{1}{\lambda}\cdot\mathcal{D}_{\text{kl}}(h((z_{i})_{i=1}^{n})\|h((z^{\prime}_{i})_{i=1}^{n}))$ (11)
with $\lambda>0$.
Rather than the standard Cross-Entropy loss, the MART loss function uses a boosted version of Cross-Entropy (BCE) and focuses on the robustness of misclassified examples. Figure 1 shows an overview of our proposed method.
We now have $2$ objective functions to be simultaneously optimized. This can be done using the Conic Scalarization (CS) method (Kasimbeyli, 2013), which is proven to produce an efficient Pareto optimal point with a choice of a reference point $a$, a preference vector $k$, and an augmentation coefficient $\gamma$:
$\min_{\theta\in\Omega}\bigg{(}\sum_{i=1}^{2}k_{i}\big{(}\mathcal{L}_{i}-a_{i}\big{)}+\gamma\sum_{i=1}^{2}(\mathcal{L}_{i}-a_{i})\bigg{)}$ (CS($k,\gamma,a$))
with $(k,\gamma)\in\left\{((k_{1},k_{2}),\gamma)\mid 0\leq\gamma<k_{i},\ i=1,2\right\}$, and $20\leq a_{i}<\mathcal{L}_{i},\ i=1,$ 2.
Our multi-objective optimization approach then provides a comprehensive framework for enhancing the performance of deep neural networks under adversarial attacks.
4 Experiments
To evaluate the effectiveness of our proposed method, we perform comprehensive experiments on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) using WideResNet34-10 (Zagoruyko, 2016).
4.1 Implementation Details
In all experiments with MOREL, we use a batch size of $8$, with parameters set to $k=(0.1,0.9)$, $a=(0,0)$, $\gamma=2\times 10^{-5}$, and $\alpha=10^{-5}$. $L_{\text{e}}$ is a single linear layer with a size of $b=128$, and we use $m=2$ heads in $M_{\text{e}}$. We apply the Stochastic Gradient Descent (SGD) optimizer with a momentum factor of $0.9$, an initial learning rate of $0.01$, which is reduced by a factor of $100$ at the $75^{th}$ and $90^{th}$ epochs, and a weight decay of $10^{-4}$. The total number of epochs is set to $100$. For other methods, we use the configurations specified by their original authors (especially, $61/\lambda=$ 6 for both TRADES and MART). All methods generate adversarial examples during training using PGD-10 (Madry et al., 2017), with the maximum $l_{\infty}-\text{norm}$ of perturbations set to $\epsilon=8/255$, using random start and step size $\epsilon/4$. Across the training epochs, we evaluate all models using PGD-20 and save the best-performing model as best. The models obtained at the end of training are referred to as last. Our method, using TRADES and MART as the loss function $\mathcal{L}_{2}$, is referred to as MOREL-T and MOREL-M, respectively. All experiments are conducted on an NVIDIA A100 80GB GPU.
For testing, we use various attack methods, including FGSM (Goodfellow et al., 2014), PGD-20, PGD-100, with a step size of $\epsilon/10$, as well as $\text{CW}_{\infty}$ attack (Carlini & Wagner, 2017), using the Python library Adversarial Robustness Toolbox (ART) (Nicolae et al., 2018). For the $\text{CW}_{\infty}$ attack, we set the maximum number of iterations to $10$, with an initial learning rate of $10^{-2}$. We use a confidence level of $1$ and initialize the constant $c$ at $15$. All attack methods are evaluated under the non-targeted setting, with adversarial perturbation strength constrained by the $l_{\infty}-\text{norm}$. We evaluate both the best and last models for each method and refer to the average performance across all considered attacks as Avg-Robust.
4.2 Evaluation and Analysis of White-box Robustness
[…]
CIFAR-10 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last TRADES 84.66 $85.43$ $60.24$ $60.08$ $55.34$ $52.40$ $54.22$ $50.04$ $44.94$ 46.45 $53.69$ $52.24$ MART $82.58$ 86.12 61.57 60.83 57.27 $52.91$ 56.36 $50.68$ 47.26 $45.85$ 55.61 $52.57$ MOREL-T 85.36 85.72 $61.05$ $60.50$ $55.49$ 54.49 $54.33$ 53.12 $45.17$ $44.62$ $54.01$ 53.18 MOREL-M 82.72 $84.57$ 62.15 62.25 57.56 56.59 56.46 55.38 47.86 47.03 56.00 55.31 CIFAR-100 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last TRADES $58.41$ $58.09$ $33.73$ $31.34$ $31.25$ $27.86$ $30.73$ $26.99$ 23.25 22.21 $29.74$ $27.10$ MART $56.46$ $58.39$ 34.42 $30.21$ 31.76 $25.37$ 31.44 $24.44$ $23.14$ $20.50$ 30.19 $25.13$ MOREL-T 58.74 58.80 $33.25$ 32.85 $30.11$ 29.78 $29.55$ 29.16 $22.80$ 22.21 $28.93$ 28.50 MOREL-M 61.61 62.25 36.73 36.06 32.81 31.96 32.08 31.10 25.72 25.38 31.83 31.13
In this section, we evaluate the adversarial robustness of our proposed MOREL method under white-box attack scenarios, where the adversary has full access to the model’s parameters and gradients.
The results, presented in Table 4.2, show that for the CIFAR-10 dataset, MOREL-T achieves the highest clean accuracy of $85.36\%$ among all methods, surpassing both TRADES and MART when considering the best models. The accuracy of its last model is $85.72\%$, placing it second after MART. Under FGSM attacks, MOREL-M shows superior performance, achieving the highest accuracy of $62.15\%$ for its best model and $62.25\%$ for its last model. Against stronger attacks like PGD-20 and PGD-100, MOREL-M maintains a competitive edge, particularly with an accuracy of $56.59\%$ for its last model, under PGD-20, which is the highest across all evaluated models. For PGD-100, MOREL-M continues to show robustness, achieving $55.38\%$ with its last model. When evaluated under $\text{CW}_{\infty}$ attacks, MOREL-M again demonstrates strong performance, achieving $47.86\%$ with its best model and $47.03\%$ with its last model, surpassing TRADES and MART. The average robustness (Avg-Robust) metric across all attack types highlights the overall resilience of MOREL-M, with scores of $56.00\%$ (best) and $55.31\%$ (last), indicating consistent robustness across all white-box attacks.
For the CIFAR-100 dataset, MOREL-M leads in clean accuracy, achieving the highest accuracy of $61.61\%$ among the best models and $62.25\%$ among the last models, outperforming both TRADES and MART. This high performance in clean accuracy did not significantly compromise its robustness, as it remained dominant over other methods even under stronger attacks. The Avg-Robust scores of MOREL-M, $28.93\%$ (best) and $28.50\%$ (last), further demonstrate its consistent resilience to white-box attacks on the CIFAR-100 dataset.
The consistent performance of MOREL-M with both the best and last models underscores that MOREL’s training paradigm produces robust models throughout the training process. These results demonstrate that strong feature representation learning, as achieved by the MOREL framework, significantly enhances adversarial robustness.
4.3 Evaluation and Analysis of black-box Robustness
[…]
CIFAR-10 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last TRADES 82.57 $83.68$ 83.24 $84.24$ 83.14 $84.01$ 84.40 $85.13$ 83.34 $84.27$ MART $80.23$ 84.31 $81.13$ 84.75 $80.93$ 84.67 $82.33$ 85.86 $81.16$ 84.90 MOREL-T 83.25 83.84 83.98 84.34 83.90 84.22 85.09 85.44 84.06 84.46 MOREL-M $80.63$ $82.42$ $81.33$ $83.19$ $81.05$ $82.97$ $82.47$ $84.32$ $81.37$ $83.23$ CIFAR-100 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last TRADES 56.43 $56.19$ 56.53 $56.25$ $56.39$ $56.12$ $58.07$ $57.80$ 56.86 $56.59$ MART $54.28$ $55.73$ $54.49$ $56.10$ $54.32$ $55.89$ $56.21$ $58.02$ $54.83$ $56.44$ MOREL-T $55.99$ 56.53 $56.49$ 56.87 56.40 56.66 58.36 58.38 $56.81$ 57.11 MOREL-M 58.82 59.63 59.30 59.92 58.98 59.54 61.22 62.02 59.58 60.28
In addition to white-box attacks, we evaluate the robustness of our models against black-box attacks, where the adversary does not have direct access to the model’s parameters or gradients. Table 4.3 presents the performance of MOREL-T and MOREL-M compared to TRADES and MART on both the CIFAR-10 and CIFAR-100 datasets.
For the CIFAR-10 dataset, MOREL-T demonstrates strong robustness against all attacks, achieving an accuracy of $85.09\%$ against the $\text{CW}_{\infty}$ attack, outperforming both TRADES and MART when considering the best models. While MART shows superior performance for the last models, MOREL-T maintains consistently high robustness across both the best and last models. In the CIFAR-100 dataset, MOREL-M excels in FGSM attacks, achieving the highest accuracy of $58.82\%$ among the best models and an accuracy of $59.63\%$ among the last models, outperforming all other methods. Under PGD-20 and PGD-100 attacks, MOREL-M continues to demonstrate strong robustness, with best accuracies of $59.30\%$ and $58.98\%$, respectively. In the $\text{CW}_{\infty}$ attack scenario, MOREL-T also performs well, achieving $58.36\%$ with its best model and $57.11\%$ with its last model. For Avg-Robust, MOREL-M leads the performance in CIFAR-100 with scores of $56.81\%$ (best) and $57.11\%$ (last), indicating strong overall robustness under black-box attacks.
The results in black-box settings further reinforce the effectiveness of our multi-objective learning framework. Particularly in the challenging CIFAR-100 dataset, MOREL-T and MOREL-M consistently outperform TRADES and MART. This indicates that our method effectively generalizes across attack types.
5 Conclusion
In this paper, we introduced MOREL, a multi-objective feature representation learning framework aimed at enhancing the adversarial robustness of deep neural networks. MOREL encourages the alignment of natural and adversarial features through the use of cosine similarity and contrastive losses during training, promoting the learning of robust feature representations. Our approach consistently outperformed existing methods that similarly require no architectural changes or test-time data purification, such as TRADES and MART, in terms of robustness against a wide range of adversarial attacks, while maintaining high clean accuracy. Moreover, the ability of our multi-objective optimization approach to generalize across various datasets and attack types, without requiring architectural modifications, makes it a practical and scalable solution for real-world applications. For future work, we plan to investigate the transferability of robust features learned by MOREL across different tasks and domains, which could unlock new possibilities for applying adversarially robust models in areas like transfer learning and domain adaptation. Additionally, we aim to explore grouping techniques to extend MOREL’s application to scenarios with limited labeled data, such as semi-supervised or few-shot learning settings.
6 Acknowledgement
This project received funding from the German Federal Ministry of Education and Research (BMBF) through the AI junior research group “Multicriteria Machine Learning”.

### v4

@@
-MOREL: Enhancing Adversarial Robustness through Multi-Objective Representation Learning
+Enhancing Adversarial Robustness through Multi-Objective Representation Learning
 Abstract
-Extensive research has shown that deep neural networks (DNNs) are vulnerable to slight adversarial perturbations—small changes to the input data that appear insignificant but cause the model to produce drastically different outputs.
-In addition to augmenting training data with adversarial examples generated from a specific attack method, most of the current defense strategies necessitate modifying the original model architecture components to improve robustness or performing test-time data purification to handle adversarial attacks.
-In this work, we demonstrate that strong feature representation learning during training can significantly enhance the original model’s robustness.
-We propose MOREL, a multi-objective feature representation learning approach, encouraging classification models to produce similar features for inputs within the same class, despite perturbations.
-Our training method involves an embedding space where cosine similarity loss and multi-positive contrastive loss are used to align natural and adversarial features from the model encoder and ensure tight clustering.
-Concurrently, the classifier is motivated to achieve accurate predictions.
-Through extensive experiments, we demonstrate that our approach significantly enhances the robustness of DNNs against white-box and black-box adversarial attacks, outperforming other methods that similarly require no architectural changes or test-time data purification.
+Deep neural networks (DNNs) are vulnerable to small adversarial perturbations, which are tiny changes to the input data that appear insignificant but cause the model to produce drastically different outputs.
+Many defense methods require modifying model architectures during evaluation or performing test-time data purification.
+This not only introduces additional complexity but is often architecture-dependent.
+We show, however, that robust feature learning during training can significantly enhance DNN robustness.
+We propose MOREL, a multi-objective approach that aligns natural and adversarial features using cosine similarity and multi-positive contrastive losses to encourage similar features for same-class inputs.
+Extensive experiments demonstrate that MOREL significantly improves robustness against both white-box and black-box attacks.
 Our code is available at https://github.com/salomonhotegni/MOREL.
+Keywords:
+Adversarial robustness Multi-objective optimization.
 1 Introduction
-The deployment of deep neural networks (DNNs) in critical vision applications such as autonomous driving and medical diagnosis (Bojarski et al., 2016; Miotto et al., 2018) underscores the need for robust models capable of reliably handling real-world scenarios.
-However, extensive research has demonstrated that DNNs are vulnerable to adversarial examples—inputs crafted by adding imperceptible perturbations that can cause the model to make incorrect predictions with high confidence (Nguyen et al., 2015; Szegedy et al., 2013; Goodfellow et al., 2014).
-This vulnerability poses significant challenges to the security and reliability of AI systems, especially in safety-critical environments.
+Deep neural networks (DNNs) have achieved impressive results in many vision tasks, but their ability to generalize beyond the training distribution remains a major challenge, especially when deployed in safety-critical domains such as autonomous driving and medical diagnosis [2, 20].
+In particular, extensive research has demonstrated that DNNs can be fooled by adversarial examples: inputs that differ only by imperceptible perturbations yet cause high-confidence misclassifications [21].
 To mitigate the risks posed by adversarial attacks, various defense strategies have been proposed.
-A common approach is adversarial training (Madry et al., 2017), where models are trained on adversarial examples generated from a specific attack method.
-In addition, to improve robustness, most existing defenses require modifications to the original model architecture (Panousis et al., 2021; Liu et al., 2024; Mohammed et al., 2024; Zhou et al., 2023), introducing additional complexity and often being architecture-dependent.
-Some approaches also involve test-time data purification (Meng & Chen, 2017; Song et al., 2017; Cohen & Giryes, 2024; Tang & Zhang, 2024), which increases latency, limiting their practical applicability.
+A common approach is adversarial training [18], where models are trained on adversarial examples generated from a specific attack method.
+In addition, to improve robustness, most existing defenses require modifications to the original model architecture during evaluation [23, 17], introducing additional complexity and often being architecture-dependent.
+Some approaches also involve test-time data purification [19, 26], which increases latency, limiting their practical applicability.
 In this paper, we propose a novel method named Multi-Objective REpresentation Learning (MOREL) that addresses these challenges by focusing on robust feature representation learning.
@@
 The core of our approach is a multi-objective optimization framework that simultaneously optimizes two key objectives: enhancing adversarial robustness and maintaining high classification accuracy.
-We achieve this by embedding natural and adversarial features into a lower-dimensional space, where cosine similarity and contrastive loss functions are applied to align and tightly cluster these features.
-The classifier is concurrently motivated to achieve accurate predictions.
-This approach ensures that the model’s learned representations are robust to adversarial perturbations while preserving the information necessary for accurate classification.
-The embedding space used during training is discarded, allowing the model to retain its original structure and computational efficiency during inference.
-This characteristic distinguishes MOREL from many existing defense strategies that either involve architectural changes or rely on additional modules during inference.
 Through extensive experiments (Sec.
-4), we demonstrate that our approach significantly enhances the robustness of DNN models against white-box and black-box adversarial attacks, outperforming existing adversarial training methods that similarly require no architectural changes or test-time data purification, in terms of the accuracy-robustness trade-off.
+4), we demonstrate that our approach significantly enhances the robustness of DNN models against white-box and black-box adversarial attacks in terms of the accuracy-robustness trade-off, outperforming existing adversarial training methods that similarly require no architectural changes during evaluation or test-time data purification.
 In summary, our key contributions are:
@@
 2.1 Adversarial Training
-Adversarial training, introduced by Madry et al.
-(2017), has emerged as one of the most effective defenses against adversarial attacks.
+Adversarial training, introduced by [18], has emerged as one of the most effective defenses against adversarial attacks.
 The core idea involves augmenting the training data with adversarial examples generated using methods like Projected Gradient Descent (PGD).
-While standard adversarial training has proven effective against known attacks, it often results in models becoming overly specialized to the specific types of adversarial examples used during training (Tsipras et al., 2018).
+While standard adversarial training has proven effective against known attacks, it often results in models becoming overly specialized to the specific types of adversarial examples used during training [28].
 To address this limitation, several variants of adversarial training have been proposed.
-Kannan et al.
-(2018) introduced Adversarial Logit Pairing (ALP), which enhances robustness by pairing logits from adversarial and clean examples during training.
+[12] introduced Adversarial Logit Pairing (ALP), which enhances robustness by pairing logits from adversarial and clean examples during training.
 Building on this, they proposed Clean Logit Pairing (CLP), which further refines the approach by focusing specifically on randomly selected clean training examples.
-Ding et al.
-(2018) advanced the field with Max-Margin Adversarial (MMA) training, which pushes decision boundaries further from data points, thereby offering enhanced robustness.
-The TRADES method by Zhang et al.
-(2019) marked a significant leap forward by explicitly balancing the trade-off between robustness and accuracy through a regularized loss function that minimizes the Kullback-Leibler divergence between predictions on natural and adversarial examples.
-This was further refined by MART Wang et al.
-(2019), which focuses on the robustness of misclassified examples, addressing vulnerabilities near the decision boundary.
+The TRADES method by [32] marked a significant leap forward by explicitly balancing the trade-off between robustness and accuracy through a regularized loss function that minimizes the Kullback-Leibler divergence between predictions on natural and adversarial examples.
+This was further refined by MART [29], which focuses on the robustness of misclassified examples, addressing vulnerabilities near the decision boundary.
 Despite these advancements, common limitations persist, including the challenge of maintaining strong robustness while achieving high accuracy on clean data.
 Building on these state-of-the-art adversarial training methods, our approach, MOREL, addresses these challenges by strengthening the robustness of DNNs through a robust feature representation learning technique.
-By considering a multi-objective optimization framework, MOREL aims to achieve the best possible trade-offs between robustness and accuracy—an aspect that, to our knowledge, has not been fully explored in previous work.
+By considering a multi-objective optimization framework, MOREL aims to achieve the best possible trade-offs between robustness and accuracy, an aspect that, to our knowledge, has not been fully explored in previous work.
 2.2 Insights from Contrastive Learning
 To enhance the learning of robust features in the context of adversarial training, our method also draws insights from recent advances in contrastive learning.
-Contrastive learning has been shown to be effective in producing robust and well-structured feature representations by encouraging similar samples to be closer in the embedding space while pushing dissimilar samples apart (Chen et al., 2020; Gidaris et al., 2018; He et al., 2020).
-Specifically, Khosla et al.
-(2020) extend the principles of contrastive learning to a supervised setting.
+Contrastive learning has been shown to be effective in producing robust and well-structured feature representations by encouraging similar samples to be closer in the embedding space while pushing dissimilar samples apart [4].
+Specifically, [14] extend the principles of contrastive learning to a supervised setting.
 This method leverages label information to group similar examples (i.e., those sharing the same class label) closer together in the feature space.
-This work informs the design of our embedding space in MOREL, where we apply a multi-positive contrastive loss function (Khosla et al., 2020; Tian et al., 2024) to align natural and adversarial features.
+This work informs the design of our embedding space in MOREL, where we apply a multi-positive contrastive loss function [14, 27] to align natural and adversarial features.
 By doing so, MOREL not only enhances robustness against adversarial attacks but also ensures that the learned features are tightly clustered and well-separated across different classes, improving both robustness and accuracy.
 3 Methods
-We consider a supervised classification problem where a DNN model $f$ parameterized by $\theta\in\Omega$ learns to map an input image $x\in\mathbb{R}^{d}$ to a target class $f(x)=y\in\{1,...,c\}$ where $c\in\mathbb{N}$.
-An adversarial example $x^{\prime}\in\mathbb{R}^{d}$ is an image obtained by adding a small perturbation of at most $\epsilon>0$ to every pixel of $x$ such that $f(x)\neq f(x^{\prime})$.
-With a given $l_{p}$-based adversarial region $\mathcal{R}_{p}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{p}\leq\epsilon\}$, the aim of adversarial training (Madry et al., 2017) is typically to approximately minimize the risk on the data distribution $\mathcal{D}$ over adversarial examples:
-$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\max_{x^{\prime}\in\mathcal{R}_{p}(x,\epsilon)}\mathcal{L}(\theta,f(x^{\prime}),y)\right]$ (1)
-where $\mathcal{L}$ is the loss function.
-The approximate solutions to the inner maximization problem are derived using a specific attack method to generate adversarial examples, while the outer minimization problem involves training on these generated examples.
-3.1 Adversarial Attacks
-To generate adversarial examples for training, we use the Projected Gradient Descent (PGD) attack (Madry et al., 2017).
-It is an iterative method that generates adversarial examples by iteratively applying small perturbations to the input.
+We consider supervised classification problems where a DNN model $f$ parameterized by $\theta\in\Omega$ learns to map an input image $x\in\mathbb{R}^{d}$ to a target class $f(x)=y\in\{1,...,c\}$ where $c\in\mathbb{N}$.
+An adversarial example $x^{\prime}\in\mathbb{R}^{d}$ is an image obtained by adding imperceptible perturbations to $x$ such that $f(x)\neq f(x^{\prime})$.
+With a given $l_{p}$-based adversarial region $\mathcal{R}_{p}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{p}\leq\epsilon\}$ and a loss function $\mathcal{L}$, the aim of adversarial training [18] is typically to approximately minimize the risk on the data distribution $\mathcal{D}$ over adversarial examples: $\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\max_{x^{\prime}\in\mathcal{R}_{p}(x,\epsilon)}\mathcal{L}(\theta,f(x^{\prime}),y)\right].$ The approximate solutions to the inner maximization problem are derived using a specific attack method to generate adversarial examples, while the outer minimization problem involves training on these generated examples.
+To generate adversarial examples for training, we use the Projected Gradient Descent (PGD) attack [18].
+PGD is a method that generates adversarial examples by iteratively applying small perturbations to the input.
 Given an input image $x$, the true label $y$, a loss function $\mathcal{L}(\theta,x,y)$, and a model parameterized by $\theta$, the PGD attack generates an adversarial example $x^{\prime}$ through the following iterative process for a predefined number of iterations:
@@
 where, ${x^{\prime}}^{i}$ is the adversarial example at the $i$-th iteration, $\eta$ the step size, $\epsilon$ the maximum perturbation allowed, and $\text{Proj}_{\mathcal{R}(x,\epsilon)}$ the projection operator that ensures the adversarial example remains within the $\epsilon$-ball centered at $x$.
-Especially, we consider the $l_{\infty}$-based adversarial region:
-$\mathcal{R}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{\infty}\leq\epsilon\}.$
+We consider the $l_{\infty}$-based adversarial region: $\mathcal{R}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{\infty}\leq\epsilon\}.$
 3.1 Multi-Objective Representation Learning
 Training a robust model often results in a decrease in test accuracy.
-The goal of adversarial robustness is then to mitigate the trade-off between accuracy and robustness, thereby enhancing the model’s performance on both natural and adversarial examples (Zhang et al., 2019; Raghunathan et al., 2020).
+The goal of adversarial robustness is then to mitigate the trade-off between accuracy and robustness, thereby enhancing the model’s performance on both natural and adversarial examples [32, 25].
 We approach this challenge as a multi-objective optimization problem.
@@
 We denote the model encoder as $g$ (typically the model without its final layer) and the classifier as $h$ (typically the final layer).
-Let =$\mathcal{B}=\big{\{}x_{i}\in\mathbb{R}^{d}\mid i\in\{1,...,n\}\big{\}}$ be a batch of $n$ natural images with classes $\big{\{}y_{i}\in\{1,...,c\}\mid i\in\{1,...,n\}\big{\}}=\mathcal{Y}$, and =$\mathcal{B}^{\prime}=\big{\{}x_{i}^{\prime}\in\mathcal{R}_{p}(x_{i},\epsilon)\mid i\in\{1,...,n\}\big{\}}$ its adversarial batch.
+Let =$\mathcal{B}=\big{\{}x_{i}\in\mathbb{R}^{d}\mid i\in\{1,...,n\}\big{\}}$ be a batch of $n$ natural images with classes $\big{\{}y_{i}\in\{1,...,c\}\mid i\in\{1,...,n\}\big{\}}=\mathcal{Y}$, and =$\mathcal{B}^{\prime}=\big{\{}x_{i}^{\prime}\in\mathcal{R}_{p}(x_{i},\epsilon)\mid x_{i}\in\mathcal{B}\big{\}}$ its adversarial batch.
 The encoder then produces features [^1] We use the matrix notation $(z_{i})_{i=1}^{n}=Z\in\mathbb{R}^{n\times o}$, where $Z$ is the concatenation of the $n$ vectors $z_{i}$, each of dimension $o$.
-In the remainder of this paper, when we use the $\bigoplus$ operation over these matrices, it refers to a concatenation operation.:
-$g(\mathcal{B})=(z_{i})_{i=1}^{n},\ \text{and}\ g(\mathcal{B}^{\prime})=(z^{\prime}_{i})_{i=1}^{n}.$ (4)
+:
+$g(\mathcal{B})=(z_{i})_{i=1}^{n}\in\mathbb{R}^{n\times o},\ \text{and}\ g(\mathcal{B}^{\prime})=(z^{\prime}_{i})_{i=1}^{n}\in\mathbb{R}^{n\times o}.\vspace{-10pt}$ (3)
 3.1.1 Embedding Space with Class-Adaptive Multi-Head Attention
-During training, we consider an embedding space that includes a linear layer $L_{\text{e}}$ to project the features from the encoder into a lower-dimensional space:
-$L_{\text{e}}((z_{i})_{i=1}^{n})=(s_{i})_{i=1}^{n},\ \text{and}\ L_{\text{e}}((z^{\prime}_{i})_{i=1}^{n})=(s^{\prime}_{i})_{i=1}^{n}.$ (5)
+During training, we consider an embedding space that includes a linear layer $L_{\text{e}}$ of size $b$ to project the features from the encoder into a lower-dimensional space:
+$L_{\text{e}}((z_{i})_{i=1}^{n})=(s_{i})_{i=1}^{n}\in\mathbb{R}^{n\times b},\ \text{and}\ L_{\text{e}}((z^{\prime}_{i})_{i=1}^{n})=(s^{\prime}_{i})_{i=1}^{n}\in\mathbb{R}^{n\times b}.\vspace{-5pt}$ (4)
 The lower-dimensional features are then grouped according to their classes:
-$(s_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}(s_{i}^{y})_{i=1}^{n_{y}},\ \text{and}\ (s^{\prime}_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}.$ (6)
-where $n_{y}$ is the number of features of class $y$ present.
+$(s_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}(s_{i}^{y})_{i=1}^{n_{y}},\ \text{and}\ (s^{\prime}_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}.\vspace{-5pt}$ (5)
+where $n_{y}$ is the number of features of class $y$ present within the batch, and $\bigoplus$ refers to a concatenation operation.
 Additionally, a class-adaptive multi-head attention module $M_{\text{e}}$ enables interaction within each lower-dimensional feature group, resulting in richer feature representations.
-This module functions similarly to the multi-head attention mechanism in the vision transformer (Dosovitskiy et al., 2020; Xiong et al., 2020), where the linearly embedded image patches can be viewed as a lower-dimensional feature group.
+This module functions similarly to the multi-head attention mechanism in the vision transformer (ViT) [6], where the linearly embedded image patches can be viewed as a lower-dimensional feature group.
 The key distinction is that our multi-head attention module operates on features from different images (instead of features from the patches of the same image), and we omit any positional embedding mechanism since the position of a feature within its lower-dimensional feature group is irrelevant in our case (otherwise, this would imply keeping track of the position of an image within its batch).
-More precisely, given a lower-dimensional feature group $(s_{i}^{y})_{i=1}^{n_{y}}\ \left(\text{or}\ ({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$, the module $M_{\text{e}}$ produces the final embedded feature group $(t_{i}^{y})_{i=1}^{n_{y}}\ \left(\text{or}\ ({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$ via Algorithm 1.
-Algorithm 1 Class-Adaptive Multi-Head Attention
-0:  A feature group $(s_{i}^{y})_{i=1}^{n_{y}}=S_{y}\in\mathbb{R}^{n_{y}\times b}$, from class $y$.
+More precisely, given a lower-dimensional feature group $(s_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n_{y}\times b}\ \left(\text{or}\ ({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$, the module $M_{\text{e}}$ produces the final embedded feature group $(t_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n_{y}\times b}\ \left(\text{or}\ ({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$ via Algorithm 1.
+This results in a total cost of $\sum_{y}O\bigl{(}n_{y}^{2}\times b\bigr{)}$, which in the worst case is $O(n^{2}\times b)$ per forward pass.
+With a moderate batch size $n$, this quadratic cost remains manageable.
+All such groups are concatenated back along the batch dimension to form:
+$T=\bigoplus\limits_{y\in\{1,...,c\}}(t_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n\times b},\ \text{and}\ T^{\prime}=\bigoplus\limits_{y\in\{1,...,c\}}({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n\times b}.$ (6)
+This approach takes advantage of the global context understanding property of the attention mechanism [6, 9] to capture dependencies and relationships across features within the same group (class).
+During model evaluation on the test set, the embedding space is discarded, keeping the original model architecture unchanged.
+Algorithm 1 Class-Adaptive Multi-Head Attention 0:  A feature group $(s_{i}^{y})_{i=1}^{n_{y}}=S_{y}\in\mathbb{R}^{n_{y}\times b}$, from class $y$.
 0:  The availability of learnable triplet weight matrices $W_{j}^{Q}$, $W_{j}^{K}$ and $W_{j}^{V}\in\mathbb{R}^{b\times b_{j}}\ (b_{j}=b/m)$ for each head $j\in\{1,...,m\}$ as well as an additional learnable weight matrix $W^{O}\in\mathbb{R}^{mb_{j}\times b}$.
-1:  for $j=1$ to $m$ do
-2:     Normalize $S_{y}$ via layer normalization.
-3:     Project $S_{y}$ through linear transformations:$Q_{j}=S_{y}W_{j}^{Q},\ K_{j}=S_{y}W_{j}^{K},\ \text{and}\ V_{j}=S_{y}W_{j}^{V}.$
-4:     Get the attention score: $A_{j}={\tt softmax}\left(\dfrac{Q_{j}K_{j}^{T}}{\sqrt{b_{j}}}\right)$
-5:     Compute the $j^{\text{th}}$ head output $O_{j}=A_{j}V_{j}$.
-6:  end for
-7:  Concatenate the outputs from the $m$ heads and project the result through a linear transformation:$O={\tt concat}(O_{1},\dots,O_{m})W^{O}$
-8:  return  $(t_{i}^{y})_{i=1}^{n_{y}}=S_{y}+O$
-This approach takes advantage of the global context understanding property of the attention mechanism (Dosovitskiy et al., 2020; Han et al., 2022) to capture dependencies and relationships across features within the same group (class).
-During model evaluation on the test set, the embedding space is discarded, keeping the original model architecture unchanged.
-Figure 1: Overview of our proposed MOREL method.
-During training, encoder $g(\cdot)$ features are projected into a lower-dimensional space using a linear layer $L_{\text{e}}$.
-A class-adaptive multi-head attention module $M_{\text{e}}$ then facilitates interaction within each of the obtained feature groups.
-The natural and adversarial embedded features are aligned using $\mathcal{L}_{cosine}$ (8), and $\mathcal{L}_{csl}$ (9) ensures tight clustering of the class-informed embedded features.
-Simultaneously, the classifier $h(\cdot)$ is encouraged to make accurate predictions.
-During evaluation, the embedding space is discarded, preserving the original model architecture.
-3.2.2 Multi-Objective Optimization
-In multi-objective optimization, the goal is to simultaneously optimize two or more conflicting objectives, which requires balancing trade-offs to find solutions that satisfy all objectives (losses) to an acceptable degree (Hotegni et al., 2024; Coello, 2007; Marler & Arora, 2004).
-We define the loss function for robustness based on the $l_{2}-$normalized batch features from the embedding space:
-$\left\|\bigoplus\limits_{y\in\{1,...,c\}}(t_{i}^{y})_{i=1}^{n_{y}}\right\|_{2}=(t_{i})_{i=1}^{n},\ \text{and}\ \left\|\bigoplus\limits_{y\in\{1,...,c\}}({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right\|_{2}=(t^{\prime}_{i})_{i=1}^{n}.$ (7)
+1:  for $j=1$ to $m$ do 2:     Normalize $S_{y}$ via layer normalization.
+3:     Project $S_{y}$ through linear transformations:$Q_{j}=S_{y}W_{j}^{Q},\ K_{j}=S_{y}W_{j}^{K},\ \text{and}\ V_{j}=S_{y}W_{j}^{V}.$ 4:     Get the attention score: $A_{j}={\tt softmax}\left(\dfrac{Q_{j}K_{j}^{T}}{\sqrt{b_{j}}}\right)$ 5:     Compute the $j^{\text{th}}$ head output $O_{j}=A_{j}V_{j}$.
+6:  end for 7:  Concatenate the outputs from the $m$ heads and project the result through a linear transformation: $O={\tt concat}(O_{1},\dots,O_{m})W^{O}$ 8:  Output: $(t_{i}^{y})_{i=1}^{n_{y}}=S_{y}+O$
+3.1.2 Multi-Objective Optimization (MOO)
+In multi-objective optimization, the goal is to simultaneously optimize two or more conflicting objectives, which requires balancing trade-offs to find solutions that satisfy all objectives (losses) to an acceptable degree [24, 11].
+We define the loss function for robustness based on $L_{\text{e}}$ outputs and the $l_{2}-$normalized batch features $T$ from the embedding space:
+$T_{\text{normalized}}=(t_{i})_{i=1}^{n}\vspace{-5pt}$ (7)
+The normalization in  (Eq. 7) computes the $l_{2}$-norm for each row (of size $b$) and divides each element in the row by this norm. This operation ensures that all feature vectors have unit norm.
 Cosine Similarity Loss:
@@
 It is calculated as follows, considering $L_{\text{e}}$ outputs:
-$\mathcal{L}_{\text{cosine}}=1-\dfrac{1}{n}\sum_{i=1}^{n}\frac{t_{i}\cdot t^{\prime}_{i}}{\|t_{i}\|\|t^{\prime}_{i}\|}$ (8)
+$\mathcal{L}_{\text{cosine}}=1-\dfrac{1}{n}\sum_{i=1}^{n}\frac{s_{i}\cdot s^{\prime}_{i}}{\|s_{i}\|\|s^{\prime}_{i}\|}\vspace{-5pt}$ (8)
 […]
+Figure 1: Overview of our proposed MOREL method.
+For each batch feature tensor $(z_{i})_{i=1}^{n}\in\mathbb{R}^{n\times o}$ (Eq.
+3) from the encoder, the linear layer $L_{\text{e}}$ produces $(s_{i})_{i=1}^{n}\in\mathbb{R}^{n\times b}$ (Eq.
+4); these are then grouped by class into: $(s_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n_{y}\times b},\ \sum_{y}n_{y}=n$ (Eq.
+5), where the group sizes $n_{y}$ may vary across $y$, while the feature dimension $b$ remains fixed.
+Each feature group is separately processed by the $M_{\text{e}}$ module, which natively handles input sequences of arbitrary length $n_{y}$ with a fixed embedding dimension $b$ (Algorithm 1).
+During evaluation, the embedding space is discarded, preserving the original model architecture.
 Multi-Positive Contrastive Loss:
-The multi-positive contrastive loss function (Khosla et al., 2020; Tian et al., 2024) encourages the model to bring the features of the same class closer while pushing the features of different classes apart, considering the natural features from $M_{\text{e}}$:
-$\mathcal{L}_{csl}=\sum_{j\in\{1,...,2n\}}\dfrac{-1}{|\mathcal{P}(j)|}\sum_{p\in\mathcal{P}(j)}\text{log}\dfrac{\text{exp}(t_{j}\cdot t_{p}/\tau)}{\sum_{q\in\mathcal{Q}(j)}\text{exp}(t_{j}\cdot t_{q}/\tau)}$ (9)
+The multi-positive contrastive loss function [14, 27] encourages the model to bring the features of the same class closer while pushing the features of different classes apart, considering the natural features $T_{\text{normalized}}$ from $M_{\text{e}}$:
+$\mathcal{L}_{csl}=\sum_{j\in\{1,...,2n\}}\dfrac{-1}{|\mathcal{P}(j)|}\sum_{p\in\mathcal{P}(j)}\text{log}\dfrac{\text{exp}(t_{j}\cdot t_{p}/\tau)}{\sum_{q\in\mathcal{Q}(j)}\text{exp}(t_{j}\cdot t_{q}/\tau)}\vspace{-5pt}$ (9)
 where $\tau\in\mathbb{R}^{+}$ is a scalar parameter, $\mathcal{Q}(j)=\{1,...,2n\}\setminus\{j\}$ and $\mathcal{P}(j)=\{p\in\mathcal{Q}(j)\mid y_{p}=y_{j}\}$ with $y_{p}$ and $y_{j}$ the class labels of $t_{p}$ and $t_{j}$.
 The loss function for robustness is then defined as follows:
-$\mathcal{L}_{1}=\mathcal{L}_{\text{cosine}}+\alpha\mathcal{L}_{csl}$ (10)
-with $10<\alpha<$ 1.
-To improve accuracy, we recommend using loss functions that induce robustness in the classifier $h(\cdot)$, such as TRADES (Zhang et al., 2019) or MART (Wang et al., 2019).
-TRADES is defined as a Cross-Entropy loss ($\mathcal{L}_{\text{ce}}$) regularized by the Kullback-Leibler divergence ($\mathcal{D}_{\text{kl}}$) between the model’s predictions on natural and adversarial examples:
-$\mathcal{L}_{2}=\mathcal{L}_{\text{ce}}(h((z_{i})_{i=1}^{n}),\mathcal{Y})+\frac{1}{\lambda}\cdot\mathcal{D}_{\text{kl}}(h((z_{i})_{i=1}^{n})\|h((z^{\prime}_{i})_{i=1}^{n}))$ (11)
-with $\lambda>0$.
-Rather than the standard Cross-Entropy loss, the MART loss function uses a boosted version of Cross-Entropy (BCE) and focuses on the robustness of misclassified examples.
-Figure 1 shows an overview of our proposed method.
-We now have $2$ objective functions to be simultaneously optimized.
-This can be done using the Conic Scalarization (CS) method (Kasimbeyli, 2013), which is proven to produce an efficient Pareto optimal point with a choice of a reference point $a$, a preference vector $k$, and an augmentation coefficient $\gamma$:
-$\min_{\theta\in\Omega}\bigg{(}\sum_{i=1}^{2}k_{i}\big{(}\mathcal{L}_{i}-a_{i}\big{)}+\gamma\sum_{i=1}^{2}(\mathcal{L}_{i}-a_{i})\bigg{)}$ (CS($k,\gamma,a$))
+$\mathcal{L}_{1}=\mathcal{L}_{\text{cosine}}+\alpha\mathcal{L}_{csl}\vspace{-5pt}$ (10)
+with $0\leq\alpha\leq 1$.
+By construction, the loss function $\mathcal{L}_{1}$ (Eq.
+10) is specifically designed to induce robustness in the encoder $g(\cdot)$.
+To enhance the model’s overall accuracy, we recommend incorporating an additional loss function $\mathcal{L}_{2}$ that promotes both accurate predictions by the classifier $h(\cdot)$ and robustness.
+We consider three baselines for the loss function $\mathcal{L}_{2}$, namely TRADES [32], MART [29], and LOAT [30].
+In the next section, when we use a “baseline” method as the loss function $\mathcal{L}_{2}$ within the MOREL framework, we refer to it as “MOREL($\leftarrow$ baseline)”.
+We now have two objective functions to be simultaneously optimized.
+This can be done using the Conic Scalarization (CS) method [13], which is proven to produce an efficient Pareto optimal point with a choice of a reference point $a$, a preference vector $k$, and an augmentation coefficient $\gamma$:
+$\min_{\theta\in\Omega}\bigg{(}\sum_{i=1}^{2}k_{i}\big{(}\mathcal{L}_{i}-a_{i}\big{)}+\gamma\sum_{i=1}^{2}(\mathcal{L}_{i}-a_{i})\bigg{)}\vspace{-5pt}$ (CS($k,\gamma,a$))
 with $(k,\gamma)\in\left\{((k_{1},k_{2}),\gamma)\mid 0\leq\gamma<k_{i},\ i=1,2\right\}$, and $20\leq a_{i}<\mathcal{L}_{i},\ i=1,$ 2.
 Our multi-objective optimization approach then provides a comprehensive framework for enhancing the performance of deep neural networks under adversarial attacks.
+Figure 1 shows an overview of our proposed method.
+[…]
+ResNet18 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ AutoAttack Avg-Robust best last best last best last best last best last best last best last CIFAR-10 TRADES 79.00 $79.41$ $53.83$ $53.74$ $49.94$ $49.31$ $49.08$ $48.60$ $39.41$ $39.03$ 69.63 $69.85$ $52.38$ $52.10$ MOREL($\leftarrow$ TRADES) 79.96 80.35 $54.72$ $54.33$ $50.64$ $49.67$ $49.84$ $48.73$ $39.65$ $39.51$ 70.76 70.76 $53.12$ $52.60$ MART $77.94$ $79.57$ $55.74$ $55.22$ $51.63$ $49.89$ $50.80$ $48.56$ $41.40$ $40.44$ $68.18$ $69.57$ $53.54$ $52.73$ MOREL($\leftarrow$ MART) $78.56$ $80.09$ 56.15 55.86 52.08 50.18 51.08 49.01 41.75 40.58 $69.18$ $69.91$ 54.05 53.10 LOAT $78.09$ $79.47$ $55.67$ $55.23$ $51.70$ $49.89$ $50.87$ $48.61$ $41.20$ $40.44$ $68.34$ $69.55$ $53.55$ $52.74$ MOREL($\leftarrow$ LOAT) $78.13$ 80.49 56.27 55.62 51.99 50.23 51.05 48.96 42.01 41.00 $68.40$ 70.45 53.94 53.25 CIFAR-100 TRADES 52.68 $52.90$ $28.41$ $28.03$ $26.21$ $25.84$ $25.90$ $25.42$ $18.21$ $18.35$ $39.96$ $40.00$ $27.73$ $27.52$ MOREL($\leftarrow$ TRADES) 56.56 55.39 $28.88$ $27.98$ $25.91$ $25.27$ $25.51$ $24.85$ $18.25$ $18.17$ 43.23 41.41 $28.35$ $27.53$ MART $51.41$ $52.40$ $28.80$ $28.22$ $26.51$ $25.25$ $26.11$ $24.76$ $18.77$ $18.14$ $39.44$ $39.44$ $27.92$ $27.16$ MOREL($\leftarrow$ MART) $52.36$ $53.26$ 30.43 29.73 28.12 27.19 27.67 26.71 20.35 19.69 $40.90$ 40.85 29.49 28.83 LOAT $51.39$ $51.90$ $28.70$ $27.57$ $25.89$ $25.18$ $25.49$ $24.61$ 19.09 $18.43$ $39.19$ $39.28$ $27.66$ $27.01$ MOREL($\leftarrow$ LOAT) $52.58$ 53.43 30.75 29.35 28.44 26.67 28.04 26.15 20.35 19.54 41.38 $40.74$ 29.78 28.49 Tiny- ImageNet TRADES 41.97 40.91 $18.91$ $18.28$ $17.31$ $16.70$ $16.99$ $16.44$ $10.06$ $09.90$ $32.02$ $30.76$ $19.06$ $18.41$ MOREL($\leftarrow$ TRADES) 43.74 42.20 $18.89$ $18.24$ $16.95$ $16.14$ $16.70$ $15.89$ $10.63$ $09.99$ 33.89 32.25 $19.41$ $18.50$ MART $39.62$ $39.90$ 21.73 $19.84$ 20.39 18.25 20.24 17.96 $12.82$ $11.58$ $31.30$ $31.14$ $21.29$ $19.75$ MOREL($\leftarrow$ MART) $40.50$ $40.89$ $21.54$ 20.73 $20.15$ 18.97 $19.92$ 18.62 13.55 12.51 $32.35$ $31.87$ 21.50 20.54 LOAT $39.33$ $39.59$ $21.23$ $19.53$ 20.24 $17.81$ $19.98$ $17.50$ $12.97$ $11.42$ $31.27$ $30.89$ $21.13$ $19.42$ MOREL($\leftarrow$ LOAT) $40.93$ $41.88$ 21.66 20.00 20.24 $17.67$ 20.02 $17.30$ 13.09 11.65 32.71 32.77 21.54 19.87
+[…]
+WideResNet34-10 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last CIFAR-10 TRADES 84.66 $85.43$ $60.24$ $60.08$ $55.34$ $52.40$ $54.22$ $50.04$ $44.94$ 46.45 $53.69$ $52.24$ MOREL($\leftarrow$ TRADES) 85.36 85.72 $61.05$ $60.50$ $55.49$ 54.49 $54.33$ 53.12 $45.17$ $44.62$ $54.01$ 53.18 MART $82.58$ 86.12 61.57 60.83 57.27 $52.91$ 56.36 $50.68$ 47.26 $45.85$ 55.61 $52.57$ MOREL($\leftarrow$ MART) $82.72$ $84.57$ 62.15 62.25 57.56 56.59 56.46 55.38 47.86 47.03 56.00 55.31 CIFAR-100 TRADES $58.41$ $58.09$ $33.73$ $31.34$ $31.25$ $27.86$ $30.73$ $26.99$ 23.25 22.21 $29.74$ $27.10$ MOREL($\leftarrow$ TRADES) 58.74 58.80 $33.25$ 32.85 $30.11$ 29.78 $29.55$ 29.16 $22.80$ 22.21 $28.93$ 28.50 MART $56.46$ $58.39$ 34.42 $30.21$ 31.76 $25.37$ 31.44 $24.44$ $23.14$ $20.50$ 30.19 $25.13$ MOREL($\leftarrow$ MART) 61.61 62.25 36.73 36.06 32.81 31.96 32.08 31.10 25.72 25.38 31.83 31.13
 4 Experiments
-To evaluate the effectiveness of our proposed method, we perform comprehensive experiments on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) using WideResNet34-10 (Zagoruyko, 2016).
 4.1 Implementation Details
-In all experiments with MOREL, we use a batch size of $8$, with parameters set to $k=(0.1,0.9)$, $a=(0,0)$, $\gamma=2\times 10^{-5}$, and $\alpha=10^{-5}$.
-$L_{\text{e}}$ is a single linear layer with a size of $b=128$, and we use $m=2$ heads in $M_{\text{e}}$.
-We apply the Stochastic Gradient Descent (SGD) optimizer with a momentum factor of $0.9$, an initial learning rate of $0.01$, which is reduced by a factor of $100$ at the $75^{th}$ and $90^{th}$ epochs, and a weight decay of $10^{-4}$.
-The total number of epochs is set to $100$.
-For other methods, we use the configurations specified by their original authors (especially, $61/\lambda=$ 6 for both TRADES and MART).
-All methods generate adversarial examples during training using PGD-10 (Madry et al., 2017), with the maximum $l_{\infty}-\text{norm}$ of perturbations set to $\epsilon=8/255$, using random start and step size $\epsilon/4$.
-Across the training epochs, we evaluate all models using PGD-20 and save the best-performing model as best.
-The models obtained at the end of training are referred to as last.
-Our method, using TRADES and MART as the loss function $\mathcal{L}_{2}$, is referred to as MOREL-T and MOREL-M, respectively.
-All experiments are conducted on an NVIDIA A100 80GB GPU.
-For testing, we use various attack methods, including FGSM (Goodfellow et al., 2014), PGD-20, PGD-100, with a step size of $\epsilon/10$, as well as $\text{CW}_{\infty}$ attack (Carlini & Wagner, 2017), using the Python library Adversarial Robustness Toolbox (ART) (Nicolae et al., 2018).
-For the $\text{CW}_{\infty}$ attack, we set the maximum number of iterations to $10$, with an initial learning rate of $10^{-2}$.
-We use a confidence level of $1$ and initialize the constant $c$ at $15$.
-All attack methods are evaluated under the non-targeted setting, with adversarial perturbation strength constrained by the $l_{\infty}-\text{norm}$.
-We evaluate both the best and last models for each method and refer to the average performance across all considered attacks as Avg-Robust.
-4.2 Evaluation and Analysis of White-box Robustness
+We evaluate our method on CIFAR-10 (10 classes), CIFAR-100[15](100 classes) and Tiny-ImageNet[16](200 classes) using ResNet18[10] and WideResNet34-10[31].
+In all experiments with MOREL, we use a batch size of 8 with hyperparameters $k=(0.1,0.9)$, $a=(0,0)$, $\gamma=2\times 10^{-5}$, and $\alpha=10^{-5}$, selected to satisfy the conditions in CS($k,\gamma,a$).
+We implement $L_{\text{e}}$ as a linear layer of size $b=128$, with $m=2$ heads in $M_{\text{e}}$, and train for 100 epochs; baselines use their original configurations.
+During training, adversarial examples are generated via PGD-10[18] with $\epsilon=8/255$, a random start, and step size $\epsilon/4$.
+Models are evaluated using PGD-20, saving the best-performing model as best and the final one as last.
+Experiments are conducted on an NVIDIA A100 80GB GPU.
+For testing, we employ FGSM[8], PGD-20, PGD-100, $\text{CW}_{\infty}$[3], AutoAttack[5], and SquareAttack[1] via the Adversarial Robustness Toolbox[22], all under a non-targeted $l_{\infty}$ constraint.
+Average performance over these attacks is reported as Avg-Robust.
 […]
-CIFAR-10 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last TRADES 84.66 $85.43$ $60.24$ $60.08$ $55.34$ $52.40$ $54.22$ $50.04$ $44.94$ 46.45 $53.69$ $52.24$ MART $82.58$ 86.12 61.57 60.83 57.27 $52.91$ 56.36 $50.68$ 47.26 $45.85$ 55.61 $52.57$ MOREL-T 85.36 85.72 $61.05$ $60.50$ $55.49$ 54.49 $54.33$ 53.12 $45.17$ $44.62$ $54.01$ 53.18 MOREL-M 82.72 $84.57$ 62.15 62.25 57.56 56.59 56.46 55.38 47.86 47.03 56.00 55.31 CIFAR-100 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last TRADES $58.41$ $58.09$ $33.73$ $31.34$ $31.25$ $27.86$ $30.73$ $26.99$ 23.25 22.21 $29.74$ $27.10$ MART $56.46$ $58.39$ 34.42 $30.21$ 31.76 $25.37$ 31.44 $24.44$ $23.14$ $20.50$ 30.19 $25.13$ MOREL-T 58.74 58.80 $33.25$ 32.85 $30.11$ 29.78 $29.55$ 29.16 $22.80$ 22.21 $28.93$ 28.50 MOREL-M 61.61 62.25 36.73 36.06 32.81 31.96 32.08 31.10 25.72 25.38 31.83 31.13
-In this section, we evaluate the adversarial robustness of our proposed MOREL method under white-box attack scenarios, where the adversary has full access to the model’s parameters and gradients.
-The results, presented in Table 4.2, show that for the CIFAR-10 dataset, MOREL-T achieves the highest clean accuracy of $85.36\%$ among all methods, surpassing both TRADES and MART when considering the best models.
-The accuracy of its last model is $85.72\%$, placing it second after MART.
-Under FGSM attacks, MOREL-M shows superior performance, achieving the highest accuracy of $62.15\%$ for its best model and $62.25\%$ for its last model.
-Against stronger attacks like PGD-20 and PGD-100, MOREL-M maintains a competitive edge, particularly with an accuracy of $56.59\%$ for its last model, under PGD-20, which is the highest across all evaluated models.
-For PGD-100, MOREL-M continues to show robustness, achieving $55.38\%$ with its last model.
-When evaluated under $\text{CW}_{\infty}$ attacks, MOREL-M again demonstrates strong performance, achieving $47.86\%$ with its best model and $47.03\%$ with its last model, surpassing TRADES and MART.
-The average robustness (Avg-Robust) metric across all attack types highlights the overall resilience of MOREL-M, with scores of $56.00\%$ (best) and $55.31\%$ (last), indicating consistent robustness across all white-box attacks.
-For the CIFAR-100 dataset, MOREL-M leads in clean accuracy, achieving the highest accuracy of $61.61\%$ among the best models and $62.25\%$ among the last models, outperforming both TRADES and MART.
-This high performance in clean accuracy did not significantly compromise its robustness, as it remained dominant over other methods even under stronger attacks.
-The Avg-Robust scores of MOREL-M, $28.93\%$ (best) and $28.50\%$ (last), further demonstrate its consistent resilience to white-box attacks on the CIFAR-100 dataset.
-The consistent performance of MOREL-M with both the best and last models underscores that MOREL’s training paradigm produces robust models throughout the training process.
-These results demonstrate that strong feature representation learning, as achieved by the MOREL framework, significantly enhances adversarial robustness.
-4.3 Evaluation and Analysis of black-box Robustness
+ResNet18 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ SquareAttack Avg-Robust best last best last best last best last best last best last CIFAR-10 TRADES 77.21 $77.61$ 77.66 $78.01$ 77.35 $77.81$ 78.75 $79.14$ 46.45 46.33 71.48 $71.77$ MOREL($\leftarrow$ TRADES) 77.84 78.59 78.27 78.88 78.16 $78.59$ 79.73 80.07 46.64 45.91 72.12 72.40 MART $76.14$ $77.75$ $76.56$ $78.17$ $76.44$ $78.01$ $77.77$ $79.24$ $46.19$ $44.85$ $70.62$ $71.60$ MOREL($\leftarrow$ MART) $76.90$ $78.28$ $77.48$ $78.83$ $77.30$ 78.61 $78.42$ $79.85$ $46.21$ $45.27$ $71.26$ $72.16$ LOAT $76.22$ $77.89$ $76.65$ $78.33$ $76.57$ $78.16$ $77.83$ $79.28$ $46.10$ $44.89$ $70.67$ $71.70$ MOREL($\leftarrow$ LOAT) $76.09$ 78.77 $76.75$ 79.13 $76.59$ 79.05 $77.87$ 80.29 $46.29$ $45.30$ $70.72$ 72.50 CIFAR-100 TRADES 50.69 $50.92$ $50.78$ $50.90$ $50.47$ $50.80$ 52.42 $52.59$ $21.89$ $21.79$ $45.25$ $45.39$ MOREL($\leftarrow$ TRADES) 53.54 52.66 53.89 52.94 53.84 52.88 56.12 54.98 $21.70$ $21.42$ 47.82 46.97 MART $49.40$ $50.51$ $49.48$ $50.86$ $49.41$ $50.56$ $51.15$ $52.17$ $22.36$ $21.66$ $44.36$ $45.15$ MOREL($\leftarrow$ MART) $50.09$ 51.39 $50.44$ $51.54$ $50.31$ $51.19$ $51.93$ $53.05$ 23.62 23.10 $45.27$ $46.05$ LOAT $49.75$ $50.35$ $49.93$ $50.41$ $49.83$ $50.32$ $51.19$ $51.62$ $22.30$ $21.71$ $44.59$ $44.87$ MOREL($\leftarrow$ LOAT) $50.61$ $51.37$ 50.86 51.64 50.70 51.49 $52.26$ 53.22 23.61 22.67 45.61 46.07 Tiny- ImageNet TRADES 40.39 $39.25$ 40.63 $39.71$ 40.67 $39.71$ 41.84 $40.76$ $12.96$ $12.81$ 35.29 $34.45$ MOREL($\leftarrow$ TRADES) 41.45 40.45 42.08 40.88 42.13 40.93 43.44 41.95 $13.13$ $12.55$ 36.44 35.35 MART $38.36$ $38.78$ $38.70$ $39.08$ $38.59$ $39.11$ $39.44$ $39.67$ 16.13 14.76 $34.24$ $34.27$ MOREL($\leftarrow$ MART) $39.32$ $39.54$ $39.65$ $39.85$ $39.59$ $39.89$ $40.33$ $40.70$ $15.60$ 15.33 $34.89$ $35.05$ LOAT $38.02$ $38.46$ $38.32$ $38.74$ $38.39$ $38.77$ $39.14$ $39.49$ $16.00$ $14.32$ $33.97$ $33.95$ MOREL($\leftarrow$ LOAT) $39.37$ 40.15 $39.75$ 40.68 $39.74$ 40.76 $40.76$ 41.67 16.16 $14.26$ $35.15$ 35.50
 […]
-CIFAR-10 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last TRADES 82.57 $83.68$ 83.24 $84.24$ 83.14 $84.01$ 84.40 $85.13$ 83.34 $84.27$ MART $80.23$ 84.31 $81.13$ 84.75 $80.93$ 84.67 $82.33$ 85.86 $81.16$ 84.90 MOREL-T 83.25 83.84 83.98 84.34 83.90 84.22 85.09 85.44 84.06 84.46 MOREL-M $80.63$ $82.42$ $81.33$ $83.19$ $81.05$ $82.97$ $82.47$ $84.32$ $81.37$ $83.23$ CIFAR-100 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last TRADES 56.43 $56.19$ 56.53 $56.25$ $56.39$ $56.12$ $58.07$ $57.80$ 56.86 $56.59$ MART $54.28$ $55.73$ $54.49$ $56.10$ $54.32$ $55.89$ $56.21$ $58.02$ $54.83$ $56.44$ MOREL-T $55.99$ 56.53 $56.49$ 56.87 56.40 56.66 58.36 58.38 $56.81$ 57.11 MOREL-M 58.82 59.63 59.30 59.92 58.98 59.54 61.22 62.02 59.58 60.28
+WideResNet34-10 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last CIFAR-10 TRADES 82.57 $83.68$ 83.24 $84.24$ 83.14 $84.01$ 84.40 $85.13$ 83.34 $84.27$ MOREL($\leftarrow$ TRADES) 83.25 83.84 83.98 84.34 83.90 84.22 85.09 85.44 84.06 84.46 MART $80.23$ 84.31 $81.13$ 84.75 $80.93$ 84.67 $82.33$ 85.86 $81.16$ 84.90 MOREL($\leftarrow$ MART) $80.63$ $82.42$ $81.33$ $83.19$ $81.05$ $82.97$ $82.47$ $84.32$ $81.37$ $83.23$ CIFAR-100 TRADES 56.43 $56.19$ 56.53 $56.25$ $56.39$ $56.12$ $58.07$ $57.80$ 56.86 $56.59$ MOREL($\leftarrow$ TRADES) $55.99$ 56.53 $56.49$ 56.87 56.40 56.66 58.36 58.38 $56.81$ 57.11 MART $54.28$ $55.73$ $54.49$ $56.10$ $54.32$ $55.89$ $56.21$ $58.02$ $54.83$ $56.44$ MOREL($\leftarrow$ MART) 58.82 59.63 59.30 59.92 58.98 59.54 61.22 62.02 59.58 60.28
+4.2 White-Box Robustness and Performance Evaluation under AutoAttack
+In this section, we evaluate the adversarial robustness of our proposed MOREL method under AutoAttack and white-box (where the adversary has full access to the model’s parameters and gradients) attack scenarios.
+With the ResNet18 architecture (Table 3.1.2), MOREL($\leftarrow$ TRADES) stands out on all three datasets, achieving superior clean accuracy with its best model.
+It also consistently outperforms TRADES with both its best and last models against AutoAttack on all three datasets, particularly across every evaluated attack on CIFAR-10.
+Notably, throughout all the considered datasets, MOREL($\leftarrow$ TRADES), MOREL($\leftarrow$ MART) and MOREL($\leftarrow$ LOAT) each consistently outperform their respective baselines by delivering superior average robust performance along with enhanced clean accuracy.
+This indicates that our defense framework is effective at preserving natural feature representations while simultaneously enhancing robustness.
+Besides ResNet18, we also evaluate the performance of TRADES, MART, and their MOREL variants under white-box attacks using the WideResNet34-10 architecture (Table 3.1.2).
+On CIFAR-10, under the PGD-100 attack MOREL($\leftarrow$ MART) consistently outperforms both MART and TRADES by more than $5\%$ with its last model; moreover under the $\text{CW}_{\infty}$ attack, it maintains its dominance with both the best and last models.
+As shown in Figure 2 for a ViT model (patch size $4$, input size $32\times 32$) evaluated under AutoAttack, our MOREL variants surpass the GAT [7] defense method on CIFAR-10.
+GAT converts the original model architecture into a Multi-Task Learning (MTL) architecture and trains it using a multi-objective optimization approach with an auxiliary task.
+Figure 2: Comparison of defense methods applied to a ViT model on CIFAR-10 under AutoAttack, including GAT [7] evaluated with three auxiliary tasks.
+SOO stands for Single Objective Optimization methods.
+4.3 Transfer-Based Black-Box Robustness and Performance Evaluation under SquareAttack
 In addition to white-box attacks, we evaluate the robustness of our models against black-box attacks, where the adversary does not have direct access to the model’s parameters or gradients.
-Table 4.3 presents the performance of MOREL-T and MOREL-M compared to TRADES and MART on both the CIFAR-10 and CIFAR-100 datasets.
-For the CIFAR-10 dataset, MOREL-T demonstrates strong robustness against all attacks, achieving an accuracy of $85.09\%$ against the $\text{CW}_{\infty}$ attack, outperforming both TRADES and MART when considering the best models.
-While MART shows superior performance for the last models, MOREL-T maintains consistently high robustness across both the best and last models.
-In the CIFAR-100 dataset, MOREL-M excels in FGSM attacks, achieving the highest accuracy of $58.82\%$ among the best models and an accuracy of $59.63\%$ among the last models, outperforming all other methods.
-Under PGD-20 and PGD-100 attacks, MOREL-M continues to demonstrate strong robustness, with best accuracies of $59.30\%$ and $58.98\%$, respectively.
-In the $\text{CW}_{\infty}$ attack scenario, MOREL-T also performs well, achieving $58.36\%$ with its best model and $57.11\%$ with its last model.
-For Avg-Robust, MOREL-M leads the performance in CIFAR-100 with scores of $56.81\%$ (best) and $57.11\%$ (last), indicating strong overall robustness under black-box attacks.
-The results in black-box settings further reinforce the effectiveness of our multi-objective learning framework.
-Particularly in the challenging CIFAR-100 dataset, MOREL-T and MOREL-M consistently outperform TRADES and MART.
-This indicates that our method effectively generalizes across attack types.
+Apart from SquareAttack, we consider transfer-based black-box attacks, where adversarial examples are generated using a surrogate ResNet50 model (trained for $200$ epochs) and then transferred to the target models.
+The surrogate model is trained on clean images using standard training [29].
+Consequently, the same attack techniques used in white-box settings are applicable here.
+For the ResNet18 architecture (Table 4.1), the MOREL variants demonstrate notable gains in robustness across all datasets, as evidenced by their enhanced Avg-Robust scores.
+Notably, MOREL($\leftarrow$ TRADES) exhibits robust black-box defense performance relative to other methods across all datasets.
+In particular, its best model on CIFAR-100 outperforms the three baselines by approximately $2.5\%$ in terms of Avg-Robust score.
+With WideResNet34-10 (Table 4.1), MOREL($\leftarrow$ TRADES) exhibits superior robust accuracy against most attacks compared to TRADES.
+On CIFAR-100, MOREL($\leftarrow$ MART)’s performance under the $\text{CW}_{\infty}$ attack is particularly notable, outperforming MART by $4\%-5\%$ in both best and last models.
+The results in black-box settings further reinforce the effectiveness of our multi-objective learning framework, indicating that our method generalizes well across different attack types.
+4.4 Ablation Study
+We explore the Pareto front by varying the values of the preference vector $k$ for the MOREL framework (MOREL($\leftarrow$ MART)) with a ResNet18 model trained on the CIFAR-10 dataset.
+Figure LABEL:fig:main_abla provides a visualization of how the loss terms and performance against PGD-20 evolve as we adjust the values of $k_{1}$ (the weight assigned to the robustness loss $\mathcal{L}_{1}$) and $k_{2}$ (the weight assigned to the accuracy loss $\mathcal{L}_{2}$).
+As the preference shifts from prioritizing robustness ($k_{1}=0.9$) to accuracy ($k_{1}=0.1$), we observe a clear trade-off between the two objectives (Figure LABEL:fig:abla_a).
+This behavior clearly illustrates the multi-objective nature of the problem, where optimizing for one objective (accuracy or robustness) leads to a trade-off with the other.
+Figure LABEL:fig:abla_b shows the relationship between robust accuracy and the values of $k_{1}$.
+As $k_{1}$ decreases towards 0.1, robust accuracy improves, reaching its peak at $k_{1}=0.3$ This emphasizes the importance of appropriately weighting the robustness loss to improve robustness.
+Table 5: Clean and robust accuracy of the model with and without the $M_{\text{e}}$ module (and the associated contrastive loss $\mathcal{L}_{csl}$).
+$M_{\text{e}}$ (and $\mathcal{L}_{csl}$) Clean 80.09 $80.00$ PGD-20 50.18 $50.03$ PGD-100 49.01 $48.85$ AutoAttack 45.27 $45.15$
+Table 6: Comparison of methods with respect to Intra/Inter Class Distance Ratio and Cosine Similarity, considering the features of all training images and their adversarial examples.
+Method $d_{\text{intra}}/d_{\text{inter}}$ ($\downarrow$) CS ($\uparrow$) MOREL ($\leftarrow$ TRADES) 0.90649 0.99942 MOREL ($\leftarrow$ MART) 0.89614 0.99815 MOREL ($\leftarrow$ LOAT) 0.89556 0.99790
+Table 4.4 compares the performance of MOREL with and without the $M_{\text{e}}$ module.
+Since its output is only used for computing the loss $\mathcal{L}_{csl}$ (Eq.
+9), removing it corresponds to setting $\alpha=0$ (Eq.
+10).
+The robust accuracy under PGD-20 and AutoAttack is higher when the $M_{\text{e}}$ module is present than when it is removed.
+Similarly, under PGD-100, the model performs better with the Me module ($49.01\%$) than without it ($48.85\%$).
+These results suggest that the $M_{\text{e}}$ module and contrastive loss $\mathcal{L}_{csl}$ contribute modestly to improving robustness, even against stronger adversarial attacks.
+Considering the outputs $T$ and $T^{\prime}$ from the MOREL embedding space (Eq.
+6), we report in Table 4.4 the average cosine similarities between the features of all training images and their adversarial examples (PGD-10).
+In addition, considering the combination of clean and adversarial features, we measure the average pairwise distance between features within the same class (intra-class) and across different classes (inter-class), reporting their ratio.
+A lower ratio indicates better class separation and tighter intra-class clustering.
+MOREL($\leftarrow$ LOAT) achieves the best Intra/Inter Distance Ratio $(0.89556)$, while MOREL($\leftarrow$ TRADES) achieves the highest cosine similarity $(0.99942)$.
+MOREL($\leftarrow$ MART) strikes a balance, achieving a strong ratio $(0.89614)$ and cosine similarity $(0.99815)$.
 5 Conclusion
-In this paper, we introduced MOREL, a multi-objective feature representation learning framework aimed at enhancing the adversarial robustness of deep neural networks.
-MOREL encourages the alignment of natural and adversarial features through the use of cosine similarity and contrastive losses during training, promoting the learning of robust feature representations.
-Our approach consistently outperformed existing methods that similarly require no architectural changes or test-time data purification, such as TRADES and MART, in terms of robustness against a wide range of adversarial attacks, while maintaining high clean accuracy.
-Moreover, the ability of our multi-objective optimization approach to generalize across various datasets and attack types, without requiring architectural modifications, makes it a practical and scalable solution for real-world applications.
-For future work, we plan to investigate the transferability of robust features learned by MOREL across different tasks and domains, which could unlock new possibilities for applying adversarially robust models in areas like transfer learning and domain adaptation.
-Additionally, we aim to explore grouping techniques to extend MOREL’s application to scenarios with limited labeled data, such as semi-supervised or few-shot learning settings.
+We introduced MOREL, a multi-objective framework that enhances adversarial robustness by aligning natural and adversarial features with cosine similarity and contrastive losses.
+MOREL outperforms methods like TRADES, MART, and LOAT in robustness and clean accuracy without needing architectural changes at test-time.
+Its generalizability across datasets and attack types makes it practical for real-world applications.
+Future work will extend MOREL to scenarios with limited labeled data.
 6 Acknowledgement
 This project received funding from the German Federal Ministry of Education and Research (BMBF) through the AI junior research group “Multicriteria Machine Learning”.
+Appendix 0.A More Ablation Studies
+0.A.1 Evaluating Multi-Objective Optimization Strategies and Dissecting the Impact of Batch Sizes on Model Robustness
+In this section, we perform an ablation study of the Multi-Objective Optimization (MOO) method used in the MOREL framework, specifically comparing Weighted Sum (WS) and Conic Scalarization (CS).
+As shown in Figure LABEL:fig:moo_abla_a, both methods exhibit a convex Pareto front.
+However, CS (black line) achieves a better balance of the loss functions.
+Figure LABEL:fig:moo_abla_b compares the robust accuracy of the models trained using WS (red line) and CS (black line) for different values of $k_{1}$, which weights the robustness objective in the multi-objective optimization process.
+For both WS and CS, the robust accuracy reaches its peak around $k_{1}=0.3$, Where CS achieves the highest improvement, while WS falls behind.
+The robust accuracy then declines as $k_{1}$ continues to increase.
+These results highlight the advantages of Conic Scalarization over the standard Weighted Sum in balancing the competing objectives of learning robust features and making accurate predictions in adversarial training, demonstrating superior empirical performance.
+In addition, we analyze the impact of varying batch sizes during training and the presence of the $M_{\text{e}}$ module (with $\mathcal{L}_{csl}$) in the embedding space on the model’s robust accuracy.
+Figure LABEL:fig:moo_abla_c illustrates the overall robust accuracy under PGD-20 attacks as a function of batch size, with values plotted for batch sizes of $8$, $32$, $128$, and $512$.
+While larger batch sizes are commonly used in contrastive learning to leverage a diverse set of negative samples, our analysis revealed a different dynamic in MOREL.
+As the batch size increases, the model’s robustness declines.
+This trend can be attributed to the differences in training paradigms.
+In standard contrastive learning [14, 4], training typically involves two distinct steps: first, the encoder is trained to cluster features in the embedding space, and then the classifier is trained on top of the frozen encoder.
+This separation allows larger batch sizes to enhance feature learning by providing a rich diversity of negative samples, with little interference from downstream classification.
+In contrast, MOREL considers a simultaneous learning approach, optimizing both feature alignment and classification objectives through multi-objective optimization.
+As these objectives can sometimes conflict, smaller batch sizes seem to focus the optimization process on a narrower subset of samples, reducing the diversity and complexity of competing gradients in each step.
+This allows the model to resolve conflicts more effectively, maintaining a better balance between the objectives.
+0.A.2 Improving MOREL Performance with Appropriate Choice of Feature Type in $\mathcal{L}_{csl}$
+While the primary experiments with MOREL were conducted using the natural features $T$ in the $\mathcal{L}_{csl}$ loss (Eq.
+9), we additionally present an ablation study considering the adversarial features $T^{\prime}$ and the combination of both natural and adversarial features ($T\bigoplus T^{\prime}$).
+Interestingly, using $T^{\prime}$ or $T\bigoplus T^{\prime}$ results in improved robustness.
+These results highlight the potential of incorporating adversarial features in $\mathcal{L}_{csl}$, either alone or in combination with natural features, to enhance the robustness of the MOREL framework, offering valuable insights for future research and applications.
+Figure 7: Performance of MOREL ($\leftarrow$ MART) with different feature types in the $\mathcal{L}_{csl}$ loss (Eq.
+9): natural ($T$), adversarial ($T^{\prime}$), and combined ($T\bigoplus T^{\prime}$).